# Reuse-Aware Compilation for Zoned Quantum Architectures Based on Neutral Atoms

### v1

@@ Abstract
Quantum computing architectures based on neutral atoms offer large scales and high-fidelity operations. They can be heterogeneous, with different zones for storage, entangling operations, and readout. Zoned architectures improve computation fidelity by shielding idling qubits in storage from side-effect noise, unlike monolithic architectures where all operations occur in a single zone. However, supporting these flexible architectures with efficient compilation remains challenging. In this paper, we propose ZAC, a scalable compiler for zoned architectures. ZAC minimizes data movement overhead between zones with qubit reuse, i.e., keeping them in the entanglement zone if an immediate entangling operation is pending. Other innovations include novel data placement and instruction scheduling strategies in ZAC, a flexible specification of zoned architectures, and an intermediate representation for zoned architectures, ZAIR. Our evaluation shows that zoned architectures equipped with ZAC achieve a 22x improvement in fidelity compared to monolithic architectures. Moreover, ZAC is shown to have a 10% fidelity gap on average compared to the ideal solution. This significant performance enhancement enables more efficient and reliable quantum circuit execution, enabling advancements in quantum algorithms and applications. ZAC is open source at https://github.com/UCLA-VAST/ZAC
@@ VII Evaluation
We implemented our proposed algorithm in Python. We employed SciPy (v1.11.0) [36] for solving minimum weight full matching and maximal cardinality matching of a bipartite graph. All experiments were conducted on an AMD EPYC 7V13 64-Core Processor at 2450 MHz and 128 GB of RAM. Our benchmark circuits are selected from QASMBench [16] with a qubit number ranging from 14 to 98 and a gate number ranging from 41 to 630. All circuits are precompiled by Qiskit (v1.0.2) with optimization level 3 [13] for preprocessing.
@@ VII-B Fidelity Model
Figure 8: Circuit fidelity comparisons across different architectures. ‘SC’ means superconducting qubits.
@@ VII-C Monolithic Architecture vs Zoned Architecture vs Superconducting Qubit
In our evaluation results, the performance gain is summarized by the geometric mean. Fig. 8 demonstrates that ZAC outperforms all baselines for every circuit. We achieve a 1.56$\times$ and 2.33$\times$ fidelity improvement compared to the superconducting Heron and grid architectures, respectively, due to better decoherence fidelity. Table II shows the fidelity breakdown and average circuit duration for the Heron architecture and ZAC, where each fidelity term is the geometric mean among all benchmarks. Although the zoned architecture based on neutral atom have longer circuit duration compared to the superconducting qubit platform due to slow operations and qubit movement, it has less decoherence errors due to the long coherence time. However, with better 2Q gate fidelity, for circuits with short circuit duration, superconducting qubit architectures may have advantages over the zoned architecture, e.g., ising_n42 has a fidelity of 0.601 with duration 2us on SC Heron, while its fidelity on the zoned architecture is 0.37 with duration 10,439us.
@@ VIII ZAC in Fault-Tolerant Quantum Computing
In fault-tolerant quantum computing (FTQC), logical qubits are encoded into physical qubits with redundancy, enabling the detection of quantum errors. For instance, the [[8,3,2]] code, as summarized in Fig. 16a, encodes 3 logical qubits using 8 physical qubits with distance 2, allowing detection of any single-qubit error [35]. To implement logical gates, physical gates are carefully designed to manipulate logical qubits while remaining within the code space. Transversal gates are broadly defined as gates that do not propagate errors within the same block, making them preferred in FTQC. The [[8,3,2]] code has two particularly useful transversal gates (see Fig. 16a): the ‘in-block gate,’ where applying physical $T^{\dagger}$ gates equates to a combination of logical CCZ, CZ, and $Z$ gates; and the ‘inter-block gate,’ where applying physical CNOTs to corresponding qubits in two blocks equates to logical CNOTs on corresponding logical qubits. However, it is proven that no quantum error-correcting code can support universal FTQC using only transversal gates [8]. Some operations require more resource-intensive methods, such as magic state distillation [5]. Distillation protocols require a large number of qubits and are still under development, especially for emerging quantum computing platforms such as neutral atoms. Thus, we focus on FTQC compilation of circuits composed of transversal gates.
@@
Fortunately, it has been shown that a class of hypercube instantaneous quantum polynomial (hIQP) circuits [3, 11] can be implemented using transversal gates and may offer quantum advantage, as sampling from these circuits could solve certain #P-hard counting problems [6]. We consider an hIQP circuit with 384 logical qubits encoded in 128 [[8,3,2]] code blocks and 448 transversal gates. In this circuit, there are layers of in-block gates interleaved with inter-block CNOT layers. The stride of the CNOTs increases by 2 in each layer, e.g., in the first layer, CNOTs connect blocks 0 and 1, 2 and 3, etc.; in the second layer, CNOTs connect blocks 0 and 2, 1 and 3, and so on. This circuit is a scaled-up version of the 48-qubit circuit from Ref. [3], likely beyond the reach of state-of-the-art classical simulation methods [22].
@@
In the FTQC context, ZAC serves two main purposes. First, it can compile physical circuits for FTQC subroutines such as syndrome extraction and logical gadgets. For the hIQP circuits on [[8,3,2]] codes, these tasks are straightforward, so we omit them here. Second, by inputting the logical circuit, ZAC can determine the movements of logical code blocks to implement the correct CNOTs between them. The physical qubits within each logical block move together. Thus, ZAC’s qubit movement strategy directly informs how logical blocks should be repositioned to execute transversal gates. In the hIQP example, we apply ZAC to the 128 code blocks and a smaller logical-level architecture. Since each code block occupies 2 rows by 4 columns, and the physical architecture includes 7 rows and 20 columns of sites in the entanglement zone, the logical-level architecture supports $\lfloor 7/2\rfloor=3$ rows and $\lfloor 20/4\rfloor=5$ columns in the entanglement zone. Ref. [3] provides a movement heuristic that extends to cases where both the number of rows and columns are powers of 2, using only $2\times 4=8$ sites out of the 15 available. In contrast, ZAC can leverage all available sites, resulting in a compiled output with 35 Rydberg stages and a physical circuit duration of 117.847 ms.
@@ Acknowledgements
This work is funded by NSF grant 2313083. The authors would like to thank Yannick Stade, Madelyn Cain, Harry Hengyun Zhou, Marcin Kalinowski, Pengyu Liu, Dolev Bluvstein, and Prof. Mikhail D. Lukin for valuable discussions.

### v2

@@ Abstract
Quantum computing architectures based on neutral atoms offer large scales and high-fidelity operations. They can be heterogeneous, with different zones for storage, entangling operations, and readout. Zoned architectures improve computation fidelity by shielding idling qubits in storage from side-effect noise, unlike monolithic architectures where all operations occur in a single zone. However, supporting these flexible architectures with efficient compilation remains challenging. In this paper, we propose ZAC, a scalable compiler for zoned architectures. ZAC minimizes data movement overhead between zones with qubit reuse, i.e., keeping them in the entanglement zone if an immediate entangling operation is pending. Other innovations include novel data placement and instruction scheduling strategies in ZAC, a flexible specification of zoned architectures, and an intermediate representation for zoned architectures, ZAIR. Our evaluation shows that zoned architectures equipped with ZAC achieve a 22x improvement in fidelity compared to monolithic architectures. Moreover, ZAC is shown to have a 10% fidelity gap on average compared to the ideal solution. This significant performance enhancement enables more efficient and reliable quantum circuit execution, enabling advancements in quantum algorithms and applications.
@@ VII Evaluation
We implemented our proposed algorithm in Python. We employed SciPy (v1.11.0) [34] for solving minimum weight full matching and maximal cardinality matching of a bipartite graph. All experiments were conducted on an AMD EPYC 7V13 64-Core Processor at 2450 MHz and 128 GB of RAM. Our benchmark circuits are selected from QASMBench [14] with a qubit number ranging from 14 to 98 and a gate number ranging from 41 to 630. All circuits are precompiled by Qiskit (v1.2.4) with optimization level 3 [11] for preprocessing.
@@ VII-B Fidelity Model
Figure 8: Circuit fidelity comparisons across different architectures. ‘SC’ means superconducting qubits. The name of the benchmark circuit is followed by the number of program qubits and the gate number. For example, qft_n18(306/324) represents the QFT circuit with 18 qubits, 306 2Q gates and 324 1Q gates.
@@ VII-C Monolithic Architecture vs Zoned Architecture vs Superconducting Qubit
In our evaluation results, the performance gain is summarized by the geometric mean. Fig. 8 demonstrates that ZAC outperforms all baselines for every circuit. We achieve a 1.56$\times$ and 2.33$\times$ fidelity improvement compared to the superconducting Heron and grid architectures, respectively, due to better decoherence fidelity. Table II shows the fidelity breakdown and average circuit duration for the Heron architecture and ZAC, where each fidelity term is the geometric mean among all benchmarks. Although the zoned architecture based on neutral atom have longer circuit duration compared to the superconducting qubit platform due to slow operations and qubit movement, it has less decoherence errors due to the long coherence time. However, with better 2Q gate fidelity, for circuits with short circuit duration, superconducting qubit architectures may have advantages over the zoned architecture, e.g., ising_n42 has a fidelity of 0.601 and 0.722 with duration 2us and 650ns on SC Heron and Grid, respectively, with while its fidelity on the zoned architecture is 0.37 with duration 10,439us.
@@ VIII ZAC in Fault-Tolerant Quantum Computing
Recent experiments in reconfigurable neutral atom arrays have demonstrated fault-tolerant quantum algorithms using tens of logical qubits [3], highlighting the need for compilation support for fault-tolerant quantum computing (FTQC).
@@
In FTQC, logical qubits are redundantly encoded into more physical qubits to enable the detection and correction of quantum errors. For instance, the [[8,3,2]] code, as summarized in Fig. 16a, encodes 3 logical qubits using 8 physical qubits with distance 2, allowing detection of any single-qubit error [33]. To implement logical gates, physical gates are carefully designed to manipulate logical qubits while remaining within the code space. Transversal gates are broadly defined as gates that do not propagate errors within the same code block, making them preferred in FTQC. In addition, reconfigurable architectures such as trapped ions and neutral atoms can directly perform transversal gates leveraging available long-range connectivity. Two example transversal gates of the [[8,3,2]] code are displayed in Fig. 16a: the ‘in-block gate,’ where applying physical $T^{\dagger}$ gates equates to a combination of logical CCZ, CZ, and $Z$ gates; and the ‘inter-block gate,’ where applying physical CNOTs to corresponding qubits in two blocks equates to logical CNOTs on corresponding logical qubits.
@@
The [[8,3,2]] codes were recently used to implement hypercube instantaneous quantum polynomial (hIQP) circuits [3, 9]. This class of circuits can be implemented using transversal gates and may offer quantum advantage, as sampling from these circuits could solve certain #P-hard counting problems [5]. We consider an hIQP circuit with 384 logical qubits encoded in 128 [[8,3,2]] code blocks and 448 transversal gates. In this circuit, there are layers of in-block gates interleaved with inter-block CNOT layers. The stride of the CNOTs increases by 2 in each layer, generating a hypercube connectivity, e.g., in the first layer, CNOTs connect blocks 0 and 1, 2 and 3, etc.; in the second layer, CNOTs connect blocks 0 and 2, 1 and 3, and so on. This circuit is a scaled-up version of the 48-qubit circuit from Ref. [3], likely beyond the reach of state-of-the-art classical simulation methods [20].
@@
In the FTQC context, ZAC serves two main purposes. First, it can compile physical qubit movements for FTQC subroutines such as syndrome extraction and logical gates. For the hIQP circuits on [[8,3,2]] codes, these tasks are straightforward, so we omit them here. Second, by inputting the logical circuit consisting of transversal gates, ZAC can determine the movements of logical code blocks to implement the correct CNOTs between them. The physical qubits within each logical block move together. Thus, ZAC’s qubit movement strategy directly informs how logical blocks should be repositioned to execute transversal gates. In the hIQP example, we apply ZAC to the 128 code blocks and a smaller logical-level architecture. Since according to the settings in [3], each code block occupies 2 rows by 4 columns, and the physical architecture includes 7 rows and 20 columns of sites in the entanglement zone, the logical-level architecture supports $\lfloor 7/2\rfloor=3$ rows and $\lfloor 20/4\rfloor=5$ columns in the entanglement zone. Ref. [3] provides a movement heuristic that extends to cases where both the number of rows and columns are powers of 2, using only $2\times 4=8$ sites out of the 15 available. In contrast, ZAC can leverage all available sites, resulting in a compiled output with 35 Rydberg stages and a physical circuit duration of 117.847 ms.
The results above demonstrate that ZAC is an effective tool for the movements required by transversal CNOTs between code blocks. However, this addresses only one of the aspects in FTQC compilation for neutral atom arrays, and further work is needed to advance toward large-scale universal computation.
@@ Acknowledgements
This work is funded by NSF grant 2313083 and NSF grant s6488. The authors thank Y. Stade, M. Cain, D. Bluvstein, P. Liu, H. Zhou, M. Kalinowski, and Prof. M. D. Lukin for valuable discussions.